The burner on an electric stove has a power output of 2.0 kW. A 610 g stainless steel tea kettle is filled with 20∘C water and placed on the already hot burner. If it takes 3.1 min for the water to reach a boil, what volume of water, in cm3, was in the kettle? Stainless steel is mostly iron, so you can assume its specific heat is that of iron.

Express your answer using two significant figures.
___cm3
Answer #1

energy = power*time=2*1000W*3.1*60sec=372000

heat gained= miron Ciron\Delta T + mw Cw\small \Delta T

=(0.610*450*(100-20))+(1000*V*4180*(100-20)) as mass=density*volume

heat lost= heat gained

372000=(0.610*450*(100-20))+(1000*V*4180*(100-20))

372000=21960+334400000V

350040=334400000V

V=1046.7 cm^3